OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)?

OpenStudy (anonymous):

So point slope form is y-y1=m(x-x1). So have you figured out the slope yet?

OpenStudy (anonymous):

Once you figure out the slope, you fill in that number with 'm'. Then you simply pick one of the points you were given and fill in the 'x1 & y1'. Then you should be able to simply the equation to get a simply y=mx+b style of an equation.

OpenStudy (lynfran):